Two balls of equal mass are made of different materials. The specific heat capacity of the first ball is denoted by A is 645 J/(kg ° C), while the specific heat capacity of the second ball is denoted by B, 240 J / (kg °C). Suppose that both balls were initially at a temperature of 21°C, and that they were then given exactly the same amount of heat. Determine what is the final temperature of ball B, if the final temperature of ball A is 74 °C? (Solution: tB=163.43 ° C)


1
Expert's answer
2021-04-13T06:36:18-0400

Find the amount of heat using data for ball A:


"Q=Am\Delta t=645·m·(74-21)=34185m."

Exactly this amount of heat is transferred to ball B:


"Q=Bm(t_2-t_1),\\\space\\
t_2=\frac{Q}{Bm}+t_1=\frac{34185m}{240m}+21=163.44º\text{C}"